# Robust $Q$-Learning for finite ambiguity sets

### v2

@@ 3. Robust Q-Learning Algorithm
  • (c)
    Due to Equation (3.1), increasing the number of probability measures $N$ linearly increases also the computation time linearly as bloating the ambiguity set by a factor $N$ corresponds to computing $N$ times as many expectations of the form $\mathbb{E}_{\mathbb{P}^{(k)}(X_{t},a_{t}(X_{t}))}[r(X_{t},a_{t}(X_{t}),X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)]$ in order to determine $k_{t}^{*}$.
4. Extensions
We acknowledge that the assumption on a finite state space and finite ambiguity sets imposed in Section 2 might be too restrictive for most real-world applications.
To address this drawback, in this section, we indicate by studying two extensions of the presented approach its flexibility and showcase how to apply our $Q$-learning framework in more general settings relevant for practical applications in less contrived settings.
For non-robust $Q$-learning ([51]), the obstacle of finite state spaces has been overcome mainly through the use of function approximation techniques, see, e.g. [5], [26], [27] and [45]. In Section 4.2 we outline a similar extension for our setting.
Moreover, in Section 4.1, we discuss how infinitely large ambiguity sets can be approximated by finite ambiguity sets to which the setting of Section 2 applies.
@@ 4.1. Approximation of infinite ambiguity sets
[…]
for which we want to compute its $Q$-value function $Q^{*}(x,a)=\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]$. Moreover, assume that $\mathcal{P}(x,a)$ fulfils [31, Assumption 2.2] so that minimizers of the corresponding value function exist (compare [31, Theorem 2.7]). Under these assumptions, we can derive the following Lemma 4.1 that allows to approximate the corresponding $Q$-value function by the $Q$-value function of an approximating sequence of finite ambiguity sets as defined in (2.1), justifying in return to apply Algorithm 1 to compute an approximation of the $Q$-value function associated to $\mathcal{P}(x,a)$.
@@
If there exists a sequence of finite ambiguity sets $\left(\mathcal{P}^{(n)}(x,a)\right)_{n\in\mathbb{N}}\subset\mathcal{P}(x,a)$ such that for all $(x,a)\in\mathcal{X}\times A$ and all $\mathbb{P}\in\mathcal{P}(x,a)$ there exists a sequence $(\mathbb{P}^{(n)})_{n\in\mathbb{N}}$ such that [^2] Note that the conditions in (4.1) are equivalent to convergence in the Wasserstein-$1$-distance, compare, e.g. [47, Definition 6.8 (i)].
@@
$\lim_{n\rightarrow\infty}Q^{(n)}(x,a)$ $:=\lim_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V^{(n)}(X_{1})\right]$
$=\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]=:Q(x,a).$
@@
fulfils [31, Assumption 2.2] and can be approximated in the sense of Lemma 4.1 by
@@
4.2. Continuous state space
Algorithm 1 assumes a finite state space and can therefore not be applied to problems formulated on a continuous state space. Howeover, if we have a continuous state space $\mathcal{X}\subset\mathbb{R}^{d}$, finite action space $A\subset\mathbb{R}^{m}$, and a finite ambiguity set
$(x,a)$ $\twoheadrightarrow\mathcal{P}(x,a):=\left\{\mathbb{P}^{(1)}(x,a),\cdots,\mathbb{P}^{(N)}(x,a)\right\},\ \ \ \ \ \ N\in\mathbb{N},$
then we can pursue a similar approach as proposed in [27] or [45] for the non-robust setting.
To this end, we parameterize the $Q$-value function by some approximating function, where a popular choice int he field of reinforcement learning are neural networks $\mathcal{X}\times\mathcal{A}\ni(x,a)\mapsto Q_{\theta}(x,a)$ parameterized by some $\theta$. The objective is then to minimize the error between $Q_{\theta}(x,a)$ and
$\min_{k=1,\dots,N}\mathbb{E}_{\mathbb{P}^{(k)}(x,a)}\left[r(x,a,X_{1})+\alpha\max_{b\in A}Q_{\theta}(X_{1},b)\right]$
on a batch of state action pairs, where the expectation is approximated via Monte-Carlo simulation.
Hence, we sample a batch of state action pairs $(x^{i},a^{i})$, $i=1,\dots,B$ for batch size $B\in\mathbb{N}$ and minimize the loss function
$\sum_{i=1}^{B}\left(Q_{\theta}(x^{i},a^{i})-\min_{k=1,\dots,N}\frac{1}{N_{\rm MC}}\sum_{j=1}^{N_{\rm MC}}\left[r(x^{i},a^{i},X^{k,j})+\alpha\max_{b\in A}Q_{\theta}(X^{k,j},b)\right]\right)^{2}$
via stochastic gradient descent, where $X^{k,j}\sim\mathbb{P}^{(k)}(x^{i},a^{i})$ for $j=1,\dots,N_{\rm MC}$.
Note that adaptions such as double Q-learning ([45]) are easily possible also in the present robust setting.

### v3

@@ 3.1. The robust Q-Learning algorithm and its convergence
3.2. Convergence rates
Analogue to results in classical $Q$-learning ([18], [52]) and in addition to the almost sure convergence guarantee from Theorem 3.1, we provide explicit non-asymptotic high probability convergence rates for the $Q$-learning algorithm presented in Algorithm 1.
Theorem 3.4.
Let $(\widetilde{\gamma}_{t})_{t\in\mathbb{N}}$, and define for all $(x,a)\in\mathcal{X}\times A$ and for all $t\in\mathbb{N}$
(3.9) $\gamma_{t}(x,a,X_{t}):=1\hskip-3.0pt\mathrm{l}_{\left\{(x,a)=\left(X_{t},~a_{t}(X_{t})\right)\right\}}\cdot\widetilde{\gamma}_{N_{t}(x,a)}$
Moreover, let $0<\alpha<1$ and let $(x_{0},\textbf{a})\in\mathcal{X}\times\mathcal{A}$ such that
  • (i)
    For any $\delta\in(0,1)$, and any $Q_{0}\in\mathbb{R}^{|\mathcal{X}|\times|A|}$ it holds with probability [^2] Here we mean w.r.t. the probability measure $\mathbb{P}_{x_{0},\textbf{a}}$. at least $1-\delta$, for all $T\in\mathbb{N}$,
    $\max_{(x,a)\in\mathcal{X}\times A}|Q_{T}(x,a)-Q^{\ast}(x,a)|\;\leq\;$ $\left(\|Q_{0}\|_{\infty}+\frac{C_{r}}{1-\alpha}\right)\frac{\prod_{i=1}^{N_{T}^{*}-1}\left(1-\widetilde{\gamma}_{i}\right)}{{1-\alpha}}$
    $+\frac{2C_{r}}{(1-\alpha)^{2}}\sqrt{2\log\Big(\frac{2|\mathcal{X}||A|}{\delta}\Big)\sum_{t=1}^{N_{T}^{*}}\left(\widetilde{\gamma}_{t}\prod_{j=t+1}^{N_{T}^{*}-1}(1-\widetilde{\gamma}_{j})\right)^{2}}$
    $+\frac{C_{r}}{(1-\alpha)^{2}}\cdot\sqrt{\frac{2\log\left((2^{|\mathcal{X}|}-2)\cdot\left(\frac{N|\mathcal{X}||A|}{\delta}\right)\right)}{N_{T}^{*}}}$
    where $N_{T}(x,a):=\sum_{s=0}^{T}\mathbf{1}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}$ is the number of visits to $(x,a)\in\mathcal{X}\times A$ up to time $T$, and $N_{T}^{*}=\min_{(x,a)\in\mathcal{X}\times A}N_{T}(x,a)$ denotes the minimal number of visits to all state action pairs until time $T$.
  • (ii)
    With the choice $\widetilde{\gamma}_{t}=\frac{1}{t+1}$ for all $t\in\mathbb{N}$, we have
    $\max_{(x,a)\in\mathcal{X}\times A}|Q_{T}(x,a)-Q^{\ast}(x,a)|\;\leq\;$ $\frac{\left(\|Q_{0}\|_{\infty}+\frac{C_{r}}{1-\alpha}\right)}{N_{T}^{*}(1-\alpha)}+\frac{2C_{r}}{(1-\alpha)^{2}}\sqrt{\frac{2\log\Big(\frac{2|\mathcal{X}||A|}{\delta}\Big)}{N_{T}^{*}}}$
    $+\frac{C_{r}}{(1-\alpha)^{2}}\cdot\sqrt{\frac{2\log\left((2^{|\mathcal{X}|}-2)\cdot\left(\frac{N|\mathcal{X}||A|}{\delta}\right)\right)}{N_{T}^{*}}}.$
Note that the error bound in Theorem 3.4 naturally decomposes into three terms. The first two terms correspond to the stochastic approximation error of the QQ–learning recursion. The third term is specific to Algorithm 1 and reflects the statistical error arising from estimating the candidate transition kernels $\mathbb{P}^{(k)}$ via empirical measures. In the special case where the transition kernels are known, this third term vanishes.
4. Approximation of other Ambiguity sets with finite ambiguity sets
In this section we show that using finite ambiguity sets - and hence our proposed $Q$-learning algorithm - provides a versatile framework for ambiguity modeling, as it cannot only be used to model ambiguity w.r.t. a finite pool of candidate models but also to approximate a wide range of other classes of infinitely large ambiguity sets.
@@ 4.1. Approximation of infinite ambiguity sets
$\begin{aligned} \mathcal{X}\times A\ni(x,a)\twoheadrightarrow\mathcal{P}(x,a)\end{aligned}\subseteq\mathcal{M}_{1}(\mathcal{X})$
for which we want to compute its $Q$-value function $Q^{*}(x,a)=\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]$. Moreover, we assume that $\mathcal{P}(x,a)$ and the reward function fulfils the following assumption.
Assumption 4.1.
The set-valued map
$\mathcal{X}\times A$ $\rightarrow\mathcal{M}_{1}(\mathcal{X})$
$(x,a)$ $\twoheadrightarrow\mathcal{P}(x,a)$
is assumed to be nonempty, compact-valued, and continuous in the weak topology.
The above assumption ensures that minimizers of the corresponding value function exist (compare [38, Theorem 2.7]). Under these assumptions, we can derive the following Lemma 4.2 that allows to approximate the corresponding $Q$-value function by the $Q$-value function of an approximating sequence of finite ambiguity sets as defined in (2.1), justifying in return to apply Algorithm 1 to compute an approximation of the $Q$-value function associated to $\mathcal{P}(x,a)$.
@@
Assume Assumption 4.1 holds true. If there exists a sequence of finite ambiguity sets $\left(\mathcal{P}^{(n)}(x,a)\right)_{n\in\mathbb{N}}\subset\mathcal{P}(x,a)$ such that for all $(x,a)\in\mathcal{X}\times A$ and all $\mathbb{P}\in\mathcal{P}(x,a)$ there exists a sequence $(\mathbb{P}^{(n)})_{n\in\mathbb{N}}$ with $\mathbb{P}^{(n)}\in\mathcal{P}^{(n)}(x,a)$ for all $n\in\mathbb{N}$ such that [^3] Note that the conditions in (4.1) are equivalent to convergence in the Wasserstein-$1$-distance, compare, e.g. [58, Definition 6.8 (i)], and Section 4.2.
@@
$\lim_{n\rightarrow\infty}{Q}^{(n)^{*}}(x,a)$ $:=\lim_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V^{(n)}(X_{1})\right]$
$=\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]=:Q^{*}(x,a).$
@@
4.2. Approximation of Wasserstein ambiguity sets
Recently, [39] introduced a $Q$-learning algorithm that incorporates model uncertainty by allowing for distributional ambiguity within a Wasserstein-ball around a reference measure. We demonstrate that this method can be approximated by designing a suitable sequence of finite ambiguity sets, and then applying Algorithm 1. Solving the robust $Q$-learning problem for these finite sets then yields an approximation of the optimal $Q$-function corresponding to Wasserstein ambiguity.
For any $q\in\mathbb{N}$ and any probability measures $\mathbb{P}_{1},\mathbb{P}_{2}\in\mathcal{M}_{1}(\mathcal{X})$, the $q$-Wasserstein distance is defined by
$W_{q}(\mathbb{P}_{1},\mathbb{P}_{2}):=\left(\inf_{\pi\in\Pi(\mathbb{P}_{1},\mathbb{P}_{2})}\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|^{q}\,\mathrm{d}\pi(x,y)\right)^{1/q},$
where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{d}$, and where $\Pi(\mathbb{P}_{1},\mathbb{P}_{2})\subset\mathcal{M}_{1}(\mathcal{X}\times\mathcal{X})$ denotes the set of all couplings of $\mathbb{P}_{1}$ and $\mathbb{P}_{2}$.
Since we work on a finite state space, each measure $\mathbb{P}_{i}$ can be represented as
$\mathbb{P}_{i}=\sum_{x\in\mathcal{X}}a_{i,x}\delta_{x},\qquad\sum_{x\in\mathcal{X}}a_{i,x}=1,\quad a_{i,x}\geq 0,\quad i=1,2,$
where $\delta_{x}$ denotes the Dirac mass at $x\in\mathcal{X}$. In this case, the Wasserstein distance admits the equivalent discrete formulation
$W_{q}(\mathbb{P}_{1},\mathbb{P}_{2})=\left(\min_{\pi_{x,y}\in\widetilde{\Pi}(\mathbb{P}_{1},\mathbb{P}_{2})}\sum_{x,y\in\mathcal{X}}\|x-y\|^{q}\,\pi_{x,y}\right)^{1/q},$
where
$\widetilde{\Pi}(\mathbb{P}_{1},\mathbb{P}_{2}):=\left\{(\pi_{x,y})_{x,y\in\mathcal{X}}\subseteq[0,1]\ \middle|\ \sum_{x^{\prime}\in\mathcal{X}}\pi_{x^{\prime},y}=a_{2,y},\ \sum_{y^{\prime}\in\mathcal{X}}\pi_{x,y^{\prime}}=a_{1,x},\ \forall\,x,y\in\mathcal{X}\right\}.$
Fix $\varepsilon>0$ and $q\in\mathbb{N}$. For each $(x,a)\in\mathcal{X}\times A$, we define the ambiguity set
(4.2) $\mathcal{P}^{(q,\varepsilon)}(x,a):=\left\{\mathbb{P}\in\mathcal{M}_{1}(\mathcal{X})\,\middle|\,W_{q}(\mathbb{P},\widehat{\mathbb{P}}(x,a))\leq\varepsilon\right\},$
i.e., the closed $q$-Wasserstein ball of radius $\varepsilon$ centered at some reference distribution $\widehat{\mathbb{P}}(x,a)\in\mathcal{M}_{1}(\mathcal{X})$. We impose the following assumption on the reference kernel.
Assumption 4.3.
The map
$\mathcal{X}\times A$ $\rightarrow\mathcal{M}_{1}(\mathcal{X})$
$(x,a)$ $\mapsto\widehat{\mathbb{P}}(x,a)$
is assumed to be continuous in the Wasserstein-$q$ topology.
Thus, for any state–action pair $(x,a)$, the set $\mathcal{P}^{(q,\varepsilon)}(x,a)$ consists of all probability measures on $\mathcal{X}$ that are within Wasserstein distance $\varepsilon$ of the reference measure $\widehat{\mathbb{P}}(x,a)$.
Lemma 4.4.
Let Assumption 4.3 hold true. Define for $n\in\mathbb{N}$ the set of probability measures on $\mathcal{X}$ with probabilities defined on an equidistant $1/n$-spaced grid via
$G^{(n)}:=\left\{\mathbb{P}\in\mathcal{M}_{1}(\mathcal{X})~\middle|~\mathbb{P}(y)=\frac{k_{y}}{n}\text{ for some }k_{y}\in\{0,1,\dots,n\}\text{ for all }y\in\mathcal{X}\right\},$
and define for all $n\in\mathbb{N}$ the ambiguity set
(4.3) $\mathcal{X}\times A\ni(x,a)\twoheadrightarrow\mathcal{P}^{(n)}(x,a):=\mathcal{P}^{(q,\varepsilon)}(x,a)\cap G^{(n)}=\left\{\mathbb{P}\in G^{(n)}~\middle|~W_{q}(\mathbb{P},\widehat{\mathbb{P}}(x,a))\leq\varepsilon\right\}.$
  • (i)
    Let $(x,a)\in\mathcal{X}\times A$, and $\mathbb{P}\in\mathcal{P}^{(q,\varepsilon)}(x,a)$. Then, for all $p\in\mathbb{N}$ and for all $n\in\mathbb{N}$ there exists some $\mathbb{P}^{(n)}\in G^{(n)}$ with
    (4.4) $W_{p}(\mathbb{P}^{(n)},\mathbb{P})\leq\frac{\max_{y,y^{\prime}\in\mathcal{X}}\|y-y^{\prime}\|\cdot|\mathcal{X}|^{1/p}}{n^{1/p}},$
    for $|\mathcal{X}|$ denoting the number of states of $\mathcal{X}$.
  • (ii)
    We define for $(x,a)\in\mathcal{X}\times A$
    $Q^{(n)^{*}}(x,a)$ $:=\inf_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V^{(n)}(X_{1})\right],$
    $Q^{*}(x,a)$ $:=\inf_{\mathbb{P}\in\mathcal{P}^{(q,\varepsilon)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right],$
    where $V^{(n)}$ denotes the value function as defined in (2.3) associated to the finite ambiguity set $\mathcal{P}^{(n)}(x,a)$ defined in (4.3). Then, there exists some $n_{0}$ such that we have for all $n\geq n_{0}$
    $\sup_{(x,a)\in\mathcal{X}\times A}|Q^{*}(x,a)-Q^{(n)^{*}}(x,a)|\leq\frac{C}{n^{1/q}},$
    for
    $C:=\frac{(L_{r}+\alpha L_{V})(\varepsilon+\max_{y,y^{\prime}\in\mathcal{X}}\|y-y^{\prime}\|\cdot|\mathcal{X}|^{1/q})}{(1-\alpha)}<\infty,$
    and where
    $L_{r}:=\max_{(x,a)\in\mathcal{X}\times\mathcal{A}}\max_{\begin{subarray}{c}y,y^{\prime}\in\mathcal{X}\\ y\neq y^{\prime}\end{subarray}}\frac{|r(x,a,y)-r(x,a,y^{\prime})|}{\|y-y^{\prime}\|}<\infty,\qquad L_{V}:=\max_{\begin{subarray}{c}y,y^{\prime}\in\mathcal{X}\\ y\neq y^{\prime}\end{subarray}}\frac{|V^{\ast}(y)-V^{\ast}(y^{\prime})|}{\|y-y^{\prime}\|}<\infty.$
Note that the above explicit approximation rates allow in particular to determine the number of finite measures needed to obtain an approximation - given a desired precision.
4.3. Approximation of parametric ambiguity sets
We now turn to parametric uncertainty, i.e., to the case where one specifies a class of potential conditional transition kernels up to the choice of the model parameter $\theta$ which is subject to ambiguity. We define for each state-action pair $(x,a)\in\mathcal{X}\times A$ the ambiguity set
(4.5) $\mathcal{P}^{\Theta}(x,a):=\left\{\widehat{\mathbb{P}}(x,a,\theta)\in\mathcal{M}_{1}(\mathcal{X})\,\middle|\,\theta\in\Theta(x,a)\right\},$
for some state-action dependent parameter set $\Theta(x,a)\subset\mathbb{R}^{d_{\theta}}$ and some parameterized probability measure $\widehat{\mathbb{P}}(x,a,\theta)\in\mathcal{M}_{1}(\mathcal{X})$.
Lemma 4.5.
Let $\Theta(x,a)\subset\mathbb{R}^{d}$ be compact for all $(x,a)\in\mathcal{X}\times A$. And assume that for all $(x,a)\in\mathcal{X}\times A$ the map $\Theta(x,a)\ni\theta\mapsto\widehat{\mathbb{P}}(x,a,\theta)$ is continuous in the Wasserstein-1 topology. Then for all $n\in\mathbb{N}$ there exists a set-valued map
$\mathcal{X}\times A\ni(x,a)\twoheadrightarrow\Theta^{(n)}(x,a)\subset\Theta(x,a)$
which contains finitely many elements, based on which we define for all $n\in\mathbb{N}$ the ambiguity set
(4.6) $\mathcal{X}\times A\ni(x,a)\twoheadrightarrow\mathcal{P}^{(n)}(x,a):=\mathcal{P}^{\Theta^{(n)}}(x,a)=\left\{\widehat{\mathbb{P}}(x,a,\theta)\in\mathcal{M}_{1}(\mathcal{X})\,\middle|\,\theta\in\Theta^{(n)}(x,a)\right\},$
such that the assumptions of Lemma 4.2 are fulfilled, and we have
$\lim_{n\rightarrow\infty}Q^{(n)^{*}}(x,a)$ $:=\lim_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V^{(n)}(X_{1})\right]$
$=\inf_{\mathbb{P}\in\mathcal{P}^{\Theta}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]=:Q^{*}(x,a),$
where $V^{(n)}$ denotes the value function as defined in (2.3) associated to the finite ambiguity set $\mathcal{P}^{(n)}(x,a)$ defined in (4.6).
@@ 4.3. Approximation of parametric ambiguity sets
can be approximated in the sense of Lemma 4.5 by
@@
Under the additional assumption of Lipschitz-continuity of the parameter $\theta$ in the Wasserstein-distance (see Section 4.2), we are able to also derive explicit approximation rates allowing in particular to determine the needed granularity of the covering radius of $\Theta^{(n)}$ for a fixed precision.
Corollary 4.7.
Let $(x,a)\in\mathcal{X}\times\mathcal{A}$, and assume that $\Theta(x,a)\subset\mathbb{R}^{d_{\theta}}$ is compact for all $(x,a)$ and that there exists a constant $L_{\theta}<\infty$ such that the parametrization is uniformly Lipschitz in $W_{1}$, i.e.
(4.7) $W_{1}\!\big(\widehat{\mathbb{P}}(x,a,\theta),\widehat{\mathbb{P}}(x,a,\theta^{\prime})\big)\leq L_{\theta}\|\theta-\theta^{\prime}\|\qquad\text{for all }(x,a)\in\mathcal{X}\times\mathcal{A},\;\theta,\theta^{\prime}\in\Theta(x,a).$
Let $\Theta^{(n)}(x,a)\subset\Theta(x,a)$ be any finite subset and define the finite ambiguity set
$\mathcal{P}^{(n)}(x,a):=\mathcal{P}^{\Theta^{(n)}}(x,a)=\{\widehat{\mathbb{P}}(x,a,\theta)\,:\,\theta\in\Theta^{(n)}(x,a)\}.$
Define the covering radius
(4.8) $\varepsilon_{n}:=\sup_{(x,a)\in\mathcal{X}\times\mathcal{A}}\;\sup_{\theta\in\Theta(x,a)}\;\inf_{\theta^{\prime}\in\Theta^{(n)}(x,a)}\|\theta-\theta^{\prime}\|.$
Then there exists $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$,
(4.9) $\sup_{(x,a)\in\mathcal{X}\times\mathcal{A}}\big|Q^{\ast}(x,a)-Q^{(n)^{*}}(x,a)\big|\;\leq\;\frac{(L_{r}+\alpha L_{V})\,L_{\theta}}{1-\alpha}\;\varepsilon_{n}$
for
$L_{r}:=\max_{(x,a)\in\mathcal{X}\times\mathcal{A}}\max_{\begin{subarray}{c}y,y^{\prime}\in\mathcal{X}\\ y\neq y^{\prime}\end{subarray}}\frac{|r(x,a,y)-r(x,a,y^{\prime})|}{\|y-y^{\prime}\|}<\infty,\qquad L_{V}:=\max_{\begin{subarray}{c}y,y^{\prime}\in\mathcal{X}\\ y\neq y^{\prime}\end{subarray}}\frac{|V(y)-V(y^{\prime})|}{\|y-y^{\prime}\|}<\infty.$
5. Continuous state spaces
In many applications of interest the state space is naturally continuous and the assumption of a finite state space imposed in Section 2 is restrictive. In this section we outline how the robust $Q$-learning framework from Section 3 extends to continuous state spaces via function approximation, in the spirit of the classical deep $Q$-learning literature ([6], [35], [36] and [56]).
5.1. Robust Markov decision problems on continuous domains
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be a compact state space equipped with its Borel $\sigma$-algebra, and let $A$ be a finite action set. We consider a bounded reward function $r:\mathcal{X}\times A\times\mathcal{X}\to\mathbb{R}$ and a discount factor $\alpha\in(0,1)$. As in Section 2, we assume an $(x,a)$-rectangular ambiguity structure; however, in contrast to Section 2 we now allow $\mathcal{X}$ to be continuous.
For each state-action pair $(x,a)\in\mathcal{X}\times\mathcal{A}$ we are given a finite set of candidate transition kernels
(5.1) $\mathcal{P}(x,a)=\Big\{\mathbb{P}^{(1)}(x,a),\dots,\mathbb{P}^{(N)}(x,a)\Big\}\subset\mathcal{M}_{1}(\mathcal{X}).$
We interpret $\mathcal{P}(x,a)$ as the set of admissible conditional distributions of the next state $X_{1}$ given $(X_{0},A_{0})=(x,a)$ and consider the corresponding robust control objective (over an infinite time horizon) under the worst-case model chosen from these kernels.
Given a bounded function $Q:\mathcal{X}\times A\to\mathbb{R}$, define the robust Bellman operator $\mathcal{H}$ by
(5.2) $(\mathcal{H}Q)(x,a):=\min_{k\in\{1,\dots,N\}}\mathbb{E}_{\mathbb{P}^{(k)}(x,a)}\Big[r(x,a,X_{1})+\alpha\max_{b\in A}Q(X_{1},b)\Big].$
A robust optimal $Q$-function $Q^{\ast}$ is then characterized (formally) as a fixed point $Q^{\ast}=\mathcal{H}Q^{\ast}$ and the robust value function satisfies $V^{\ast}(x)=\max_{a\in A}Q^{\ast}(x,a)$.
Remark 5.1.
Under standard boundedness assumptions (see [9, Chapter 7]), $\mathcal{H}$ is a contraction w.r.t.  $\|\cdot\|_{\infty}$ with modulus $\alpha$ (the proof is identical to the finite-state case), which motivates iterative methods based on approximating $\mathcal{T}$.
5.2. Function approximation and a robust deep $Q$-learning objective
When $\mathcal{X}$ is continuous, tabular representations of $Q$ are infeasible and we employ a parametric function class
$Q_{\theta}:\mathcal{X}\times A\to\mathbb{R},\qquad\theta\in\mathbb{R}^{p},$
for instance a neural network ([28], [30]). The goal is to fit $Q_{\theta}$ such that it approximately satisfies the robust Bellman equation (5.2).
Similarly to Algorithm 1 (finite-state case), we assume that we can draw samples from each candidate kernel $\mathbb{P}^{(k)}(x,a)$, i.e., we have access to $N$ simulators that, given $(x,a)$, produce independent samples $X_{1}^{(k)}\sim\mathbb{P}^{(k)}(x,a)$.
Let $\theta^{-}$ denote a (possibly delayed) target parameter. For a given $(x,a)$ define the robust TD target
(5.3) $Y_{\theta^{-}}(x,a):=\min_{k\in\{1,\dots,N\}}\mathbb{E}_{X_{1}\sim\mathbb{P}^{(k)}(x,a)}\Big[r(x,a,X_{1})+\alpha\max_{b\in\mathcal{A}}Q_{\theta^{-}}(X_{1},b)\Big].$
In practice, the inner expectation is approximated via Monte-Carlo: for a chosen $M\in\mathbb{N}$,
(5.4) $\widehat{Y}_{\theta^{-}}(x,a):=\min_{k\in\{1,\dots,N\}}\frac{1}{M}\sum_{j=1}^{M}\Big[r(x,a,X_{1,j}^{(k)})+\alpha\max_{b\in\mathcal{A}}Q_{\theta^{-}}(X_{1,j}^{(k)},b)\Big],\qquad X_{1,j}^{(k)}\sim P^{(k)}(x,a).$
Let $\mu$ be a sampling distribution on $\mathcal{X}\times\mathcal{A}$ (e.g. induced by an exploration policy with experience replay ([19])). We propose to minimize the squared Bellman residual:
[…]
Stochastic gradient descent on (5.5) yields a robust analogue of deep $Q$-learning (DQN) for finite ambiguity sets, similar to the approach presented in [30] for Wasserstein ambiguity set.
Remark 5.2 (On the implementation).
  • (a)
    As in standard DQN ([36], [56]), using a slowly updated target parameter $\theta^{-}$ stabilizes training; the robust modification affects only the target computation through the $\min_{k=1,\dots,N}$ in (5.4).
  • (b)
    One may reduce overestimation bias by replacing the target $\max_{b\in\mathcal{A}}Q_{\theta^{-}}(X_{1,j}^{(k)},b)$ with a Double-DQN target, i.e. selecting the maximizing action using the online network and evaluating it using the target network, see [56].
  • (c)
    Relative to non-robust DQN ([36], [56]), the additional cost is linear in $N$, since for each training pair $(x,a)$ one must evaluate NN Monte-Carlo averages and take their minimum.
  • (d)
    If transitions are collected from the real environment under a single unknown kernel, one may still evaluate (5.4) by resimulating next states $X_{1,j}^{(k)}$ using the available simulators for each candidate model $k$.
Section 5.2 provides a practical extension of Algorithm 1 to continuous state spaces by replacing the tabular update with a parametric approximation and by minimizing the robust Bellman residual. Establishing global convergence guarantees for neural-network function approximation is beyond the scope of this paper; nonetheless, the robust modification is structurally minimal and preserves the key feature of our approach: only sampling access to the candidate kernels $\mathbb{P}^{(k)}(x,a)$ is required.